# Generalized Linear Bandits: Almost Optimal Regret with One-Pass Update

### v1

@@ 3.2 Jointly Efficient Method
where $\lesssim$ is used to hide constant independence of $d$, $\kappa$, $S$, $R$ and $T$. The total time complexity of Algorithm 1 is $\mathcal{O}(d^{3}+d^{2}|\mathcal{X}_{t}|)$ per round, which is independent of $T$.
Theorem 2 shows that our method achieves a regret bound of $\widetilde{\mathcal{O}}(d\sqrt{T/\kappa_{*}})$ with only $\mathcal{O}(1)$ computation cost per round. This largely improves upon the $\widetilde{\mathcal{O}}(\kappa d\sqrt{T})$ bound of Filippi et al. (2010), while enjoying a one-pass update scheme with low computational overhead. Interestingly, the result suggests that, once the local curvature of the link function is properly taken into account, self-concordant GLBs are no more difficult than linear bandits, as our method achieves comparable regret and computational efficiency to OFUL for linear bandits (Abbasi-Yadkori et al., 2011). Since logistic bandits are a special case of GLBs, Theorem 2 also advances state-of-the-art results of logistic bandits by either reducing the $\mathcal{O}(\log t)$ time complexity of Faury et al. (2022) to $\mathcal{O}(1)$, or achieving an $\mathcal{O}(\sqrt{\log T})$ improvement in the regret bound over Zhang and Sugiyama (2023).
Comparison with Sawarni et al. (2024).
@@
We note that Sawarni et al. (2024) also pursued computational efficiency, but their approach is conceptually orthogonal to ours. They reduce the computation cost by employing a rare-update strategy that limits the frequency of parameter updates. However, their method remains MLE-based and requires storing all historical data, resulting in a memory cost of $\mathcal{O}(t)$. Moreover, although the rare-update strategy yields an amortized per-round time complexity of $\mathcal{O}((\log t)^{2})$ over $T$ rounds, it still incurs a worst-case time complexity of $\mathcal{O}(t)$ in certain rounds. In contrast, our method performs a one-pass update at every iteration with $\mathcal{O}(1)$ time and space complexities per round. A promising direction for future work is to combine online estimation with rare-update techniques to further accelerate the algorithm.
Discussion with Lee and Oh (2025).
@@
The concurrent work (Lee and Oh, 2025) builds on the framework of Zhang and Sugiyama (2023) and also reports an $\mathcal{O}(\sqrt{\log T})$ improvement in multinomial logit (MNL) bandits, a different problem setting that nonetheless shares certain technical connections with the GLB problem. Notably, their technique could potentially be adapted to logistic bandits to achieve an $\mathcal{O}(\log T\sqrt{T/\kappa_{*}})$ bound with $\mathcal{O}(1)$ cost. However, we identify potential technical issues in their analysis. The argument crucially relies on a condition for the normalization factor of the truncated Gaussian distribution (see Eq.(C.15) of their paper, as also restated in (43) of appendix), an assumption that is unlikely to hold in general. A detailed discussion is provided in Appendix E.
@@ 3.3 More Discussions and Limitations
In the following, we discuss the tightness of our bound by comparing it with state-of-the-art results, as well as potential directions for further improvement.
Dependence on $T$, $\kappa$ and $d$.
For the dominant term, our regret bound matches the best-known results for GLBs using the MLE (Lee et al., 2024; Sawarni et al., 2024), with respect to its dependence on $T$, $\kappa$, and $d$. In terms of the non-leading term, Sawarni et al. (2024) achieved a slightly tighter bound, as it scales with $\kappa_{\mathcal{X}}=1/\inf_{\mathbf{x}\in\cup_{t=1}^{T}\mathcal{X}_{t}}\mu^{\prime}(\mathbf{x}^{\top}\theta_{*})$, a quantity that can be smaller than $\kappa$. In the logistic bandit case, the non-leading term is further improved to be geometry-aware, adapting more precisely to the structure of the action set (Abeille et al., 2021). We conjecture that similar improvements in the non-leading term, matching the $\kappa_{\mathcal{X}}$ dependence, might be achievable by incorporating a warm-up strategy, such as Procedure 1 in (Faury et al., 2022) or Algorithm 2 in Sawarni et al. (2024). These methods allow shifting the curvature dependence from $\mu^{\prime}(X_{t}^{\top}\theta_{t})$ to $\mu^{\prime}(X_{t}^{\top}\theta_{*})$ at only an additional constant cost, thereby attaining the refined $\kappa_{\mathcal{X}}$ scaling. However, it remains an open problem whether a geometry-aware bound, similar to that achieved in the logistic bandit setting (Abeille et al., 2021), can be obtained in the GLB setting without using MLE.
Dependence on $S$ and $R$.
The MLE-based method completely remove the dependence on $S$ and $R$ in the leading term (Lee et al., 2024), whereas our method still exhibits an $S^{2}$-dependence due to the requirement of one-pass updates. For MNL bandits, Lee and Oh (2025) showed that one can achieved improved dependence on $S$ by incorporating an adaptive warm-up procedure. It may be possible to extend their warm-up technique to GLBs for a similar improvement on $S$.
@@ 4.1 A General Recipe for OMD
To prove Theorem 1, it suffices to show that the estimation error of our OMD estimator $\lVert\theta_{t+1}-\theta_{*}\rVert_{H_{t}}$ is less than $\beta_{t}(\delta)$. The analysis begins with the following lemma, which is commonly used for the convergence or regret analysis for the OMD-type update (Chen and Teboulle, 1993; Orabona, 2019; Zhao et al., 2024). Here, we show it can serve as a general recipe for analyzing the estimation error of the OMD estimator.
@@ 4.2 Analysis for the Inverse Regret
Main Challenge. The main challenge and technical contribution of this paper lies in upper bounding the inverse regret term. Although previous works (Faury et al., 2022; Zhang and Sugiyama, 2023) have provided valuable insights, their techniques are challenging to extend to the GLB setting and remain suboptimal even for logistic bandits. The main technical difficulty in bounding the inverse regret is that $\theta_{s+1}$ is itself a function of the past losses $\ell_{s}$, which prevents the direct application of standard martingale concentration inequalities. A common strategy in prior work is to introduce an intermediate term by virtually running a full-information online learning algorithm, whose estimator $\widetilde{\theta}_{s}$ only depends on information up to time $s-1$, allowing the inverse regret to be decomposed as
@@
Our Solution.
Instead of introducing an intermediate term via virtually running an online algorithm, we propose an alternative decomposition based on the mix loss, which is defined as
@@
where $P_{s}$ is a probability distribution over $\mathbb{R}^{d}$ whose specific form will be chosen later. In general, multiple choices of $P_{s}$ are possible, and we select the one that best fits our algorithmic design. The mix loss has played a central role in the theoretical analysis of exponentially weighted methods in full-information online learning (Vovk, 2001; van der Hoeven et al., 2018). In our analysis, this term is similarly instrumental in bridging the cumulative loss gap between the underlying parameter $\theta_{*}$ and the OMD estimator $\theta_{s}$, thereby enabling a sharp bound for the estimation error. We have the following lemmas to upper bound the inverse regret based on the decomposition with mix loss:
@@
Let $\{P_{s}\}_{s=1}^{\infty}$ be a stochastic process such that $P_{s}$ is a distribution over $\mathbb{R}^{d}$ and only relies on information collected until time $s-1$. Then, for any $\delta\in(0,1]$, we have
@@
Lemma 2 shows that term (a) in (9) can be upper bounded by $\log(1/\delta)$ by choosing the mix loss as an intermediate term. The upper bound is significantly smaller than the $\mathcal{O}((\log t)^{3})$ bound in Zhang and Sugiyama (2023, Lemma 12), where the online logistic regret method by Foster et al. (2018) is employed. Notably, Lemma 2 provides the flexibility to freely choose $P_{s}$, which allows us to tailor $P_{s}$ to better approximate the behavior f the OMD estimator. The formal statement and complete proof are provided as Lemma 5 in Appendix A.
@@
$\sum_{s=1}^{t}m_{s}(P_{s})-\sum_{s=1}^{t}\ell_{s}(\theta_{s+1})\leq{}$ $\frac{1}{3\eta}\sum_{s=1}^{t}\lVert\theta_{s+1}-\theta_{s}\rVert^{2}_{H_{s}}+\frac{3}{2}d\eta\ln\bigg{(}2+\frac{2C_{\mu}t}{\lambda g(\tau)}\bigg{)},$
@@
Finally, by combining Lemma 2 and Lemma 3 with the decomposition in (28), we obtain a bound on the inverse regret term in (8), thereby completing the proof.
@@ 5 Experiment
This section evaluates the proposed method on two representative GLB problems: logistic bandits ($\mu(z)=1/(1+e^{-z})$) with bounded rewards, and Poisson bandits ($\mu(z)=e^{z}$), which pose a distinct challenge as an unbounded GLB setting. Additionally, we conducted experiments using real data from the Covertype dataset (Blackard, 1998), with more detailed results provided in Appendix F.
Compared Methods.
We compare our method against four existing algorithms for the GLB problem, including the GLM-UCB (Filippi et al., 2010), GLOC (Jun et al., 2017), RS-GLinCB (Sawarni et al., 2024), and OFUGLB (Lee et al., 2024). For the logistic bandit setting, we additionally include two algorithms specifically designed for this case: OFULog-r (Abeille et al., 2021), an MLE-based method with nearly optimal guarantees and the jointly efficient method ECOLog (Faury et al., 2022). More detailed configurations of the baselines are provided in Appendix F. All experiments are conducted over 10 independent trials, and we report the average regret and running time.
@@
Results on Logistic Bandits.
For the logistic bandits, we conduct experiments under different configurations of $S$. Specifically, the underlying parameter $\theta_{*}$ is sampled from a $d$-dimensional sphere with radius $S=\{3,5,7\}$, corresponding to $\kappa=21$, $137$, and $963$, respectively. The experimental results are shown in Figure 1. Among all methods, GLOC is the fastest but suffers from relatively large regret. OFUGLB achieves the lowest regret due to its improved dependence on both $\kappa$ and $S$, but as an MLE-based method, it incurs the highest computation cost. Our method strikes a favorable balance between statistical performance and computational efficiency. Compared to OFUGLB, it achieves substantial computational savings while incurring only a modest degradation in performance. In comparison with other jointly efficient methods, namely ECOLog for logistic bandits and RS-GLinCB for GLB, our methods achieves comparable and even slightly better performance with improved computation cost as shown in Figure 1(c). Besides, our method maintains a constant per-round cost across all regimes of $\kappa$. In contrast, the computation cost of RS-GLinCB increases with $\kappa$, as its rare-update strategy results in an update frequency that scales with $\kappa$.
Results on Poisson Bandits.
For Poisson bandits, we set the norm of the true parameter as $S\in\{3,5,7\}$. The corresponding $\kappa$ is around $18$, $127$ and $882$. Poisson bandits have unbounded rewards, whereas GLM-UCB and RS-GLinCB require a predefined upper bound on the maximum reward as a parameter. For empirical validation, we set this parameter to 100, assuming the rewards are effectively bounded with high probability. Figure 2 presents the results. Our method reduces the computation cost of the state-of-the-art OFUGLB by 1000 times, while incurring only a modest increase in regret.
@@ 6 Conclusion
This paper proposed a new method for the GLB problem that achieves a nearly optimal regret bound of $\mathcal{O}(\log T\sqrt{T/\kappa_{*}})$ with $\mathcal{O}(1)$ time and space complexities per round. Our approach builds on a novel analysis of the OMD estimator using the mix loss, enabling the construction of a tight confidence set for action selection. A natural extension of our method is to incorporate a warm-up strategy, as done in prior work, to improve the dependence on $S$ and obtain $\kappa_{\mathcal{X}}$-based bounds. Besides, it remains open whether one can achieve the geometry-aware bounds similar in GLBs to those in the logistic bandit (Abeille et al., 2021). Another direction is to relax the self-concordance assumption (Assumption 3) toward the relaxed conditions studied by Liu et al. (2024).

### v2

@@ 3.2 Jointly Efficient Method
where $\lesssim$ is used to hide constant independence of $d$, $\kappa$, $S$, $R$ and $T$.
Theorem 2 shows that our method achieves an $\widetilde{\mathcal{O}}(d\sqrt{T/\kappa_{*}})$ regret, greatly improving upon the $\widetilde{\mathcal{O}}(\kappa d\sqrt{T})$ bound of Filippi et al. (2010). From a computational perspective, our OMD estimator can be updated in $\mathcal{O}(1)$ time and memory per round, in the same spirit as the least-squares estimator in linear bandits. Consequently, the overall per-round computation cost of our algorithm matches that of LinUCB (Abbasi-Yadkori et al., 2011), indicating that the nonlinearity of the link function does not necessarily make GLBs more computationally demanding. In the finite-arm setting, our algorithm enjoys a constant per-round computational cost of $\mathcal{O}(d^{3}+d^{2}|\mathcal{X}_{t}|)$, independent of $T$. In the infinite-arm case, the arm-selection step (6) could become the main computational bottleneck (once the computational issue of MLE is resolved). Our estimator remains broadly useful as a plug-in component that can be integrated into a wide range of strategies beyond the UCB-based exploration, e.g., it can also be incorporated into Thompson Sampling (Faury et al., 2022, Appendix D.2) where the arm-selection step reduces to a convex optimization problem for convex arm sets.
Since logistic bandits are a special case of GLBs, Theorem 2 also advances state-of-the-art results of logistic bandits, either reducing the $\mathcal{O}(\log t)$ time complexity of Faury et al. (2022) to $\mathcal{O}(1)$, or achieving an $\mathcal{O}(\sqrt{\log T})$ improvement in the regret bound over Zhang and Sugiyama (2023).
@@
Comparison with Sawarni et al. (2024). We note that Sawarni et al. (2024) also pursued computational efficiency, but their approach is conceptually orthogonal to ours. They reduce the computation cost by employing a rare-update strategy that limits the frequency of parameter updates. However, their method remains MLE-based and requires storing all historical data, resulting in a memory cost of $\mathcal{O}(t)$. Moreover, although the rare-update strategy yields an amortized per-round time complexity of $\mathcal{O}((\log t)^{2})$ over $T$ rounds, it still incurs a worst-case time complexity of $\mathcal{O}(t)$ in certain rounds. In contrast, our method performs a one-pass update at every iteration with $\mathcal{O}(1)$ time and space complexities per round. A promising direction for future work is to combine online estimation with rare-update techniques to further accelerate the algorithm.
@@
Discussion with Lee and Oh (2025). The concurrent work (Lee and Oh, 2025) builds on the framework of Zhang and Sugiyama (2023) and also reports an $\mathcal{O}(\sqrt{\log T})$ improvement in multinomial logit (MNL) bandits, a different problem setting that nonetheless shares certain technical connections with GLB. Notably, their technique could potentially be adapted to logistic bandits to achieve an $\mathcal{O}(\log T\sqrt{T/\kappa_{*}})$ bound with $\mathcal{O}(1)$ cost. However, we identify potential technical issues in their analysis. The argument crucially relies on a condition for the normalization factor of the truncated Gaussian distribution (see Eq.(C.15) of their paper, as also restated in (43) of appendix), an assumption that is unlikely to hold in general. A detailed discussion is provided in Appendix D.
@@ 3.3 More Discussions and Limitations
Dependence on $T$, $\kappa$ and $d$. For the dominant term, our regret bound matches the best-known results for GLBs using the MLE (Lee et al., 2024; Sawarni et al., 2024), with respect to its dependence on $T$, $\kappa$, and $d$. In terms of the non-leading term, Sawarni et al. (2024) achieved a slightly tighter bound, as it scales with $\kappa_{\mathcal{X}}=1/\inf_{\mathbf{x}\in\cup_{t=1}^{T}\mathcal{X}_{t}}\mu^{\prime}(\mathbf{x}^{\top}\theta_{*})$, a quantity that can be smaller than $\kappa$. In the logistic bandit case, the non-leading term is further improved to be geometry-aware, adapting more precisely to the structure of the action set (Abeille et al., 2021). We conjecture that similar improvements in the non-leading term, matching the $\kappa_{\mathcal{X}}$ dependence, might be achievable by incorporating a warm-up strategy, such as Procedure 1 in (Faury et al., 2022) or Algorithm 2 in Sawarni et al. (2024). These methods allow shifting the curvature dependence from $\mu^{\prime}(X_{t}^{\top}\theta_{t})$ to $\mu^{\prime}(X_{t}^{\top}\theta_{*})$ at only an additional constant cost, thereby attaining the refined $\kappa_{\mathcal{X}}$ scaling. However, it remains an open problem whether a geometry-aware bound, similar to that achieved in the logistic bandit setting (Abeille et al., 2021), can be obtained in the GLB setting without using MLE.
Dependence on $S$ and $R$. The MLE-based method completely remove the dependence on $S$ and $R$ in the leading term (Lee et al., 2024), whereas our method still exhibits an $S^{2}$-dependence due to the requirement of one-pass updates. For MNL bandits, Lee and Oh (2025) showed that one can achieved improved dependence on $S$ by incorporating an adaptive warm-up procedure. It may be possible to extend their warm-up technique to GLBs for a similar improvement on $S$.
@@ 4.1 A General Recipe for OMD
To prove Theorem 1, it suffices to show that the estimation error $\lVert\theta_{t+1}-\theta_{*}\rVert_{H_{t}}\leq\beta_{t}(\delta)$. The analysis begins with the following lemma, which is commonly used for the convergence or regret analysis for the OMD-type update (Chen and Teboulle, 1993; Orabona, 2019; Zhao et al., 2024). Here, we show it can serve as a general recipe for analyzing the estimation error of the OMD estimator.
@@ 4.2 Analysis for the Inverse Regret
Main Challenge. The main challenge and technical contribution of this paper lies in upper bounding the inverse regret term. Although previous works (Faury et al., 2022; Zhang and Sugiyama, 2023) have provided valuable insights, their techniques are challenging to extend to the GLB setting and remain suboptimal even for logistic bandits. The main technical difficulty in bounding the inverse regret is that $\theta_{s+1}$ is itself a function of the past losses $\ell_{s}$, which prevents the direct application of standard martingale concentration inequalities. A common strategy in prior work is to introduce an intermediate term by virtually running a full-information online learning algorithm, whose estimator $\widetilde{\theta}_{s}$ only depends on information up to time s1s-1, allowing the inverse regret to be decomposed as
@@
Our Solution. Instead of introducing an intermediate term via virtually running an online algorithm, we propose an alternative decomposition based on the mix loss, which is defined as
@@
where $P_{s}$ is a probability distribution over $\mathbb{R}^{d}$ whose specific form will be chosen later. In general, several choices of $P_{s}$ are possible, and we select the one that best fits our algorithmic design. The mix loss has played a central role in the analysis of exponentially weighted methods in full-information online learning (Vovk, 2001; van der Hoeven et al., 2018). In our analysis, the mix loss is instrumental in analyzing the inverse regret defined in (9). In particular, the decomposition based on the mix loss $m_{s}(P_{s})$ offers a more general and analytically versatile formulation than $\ell_{s}(\widetilde{\theta}_{s})$ used in (9). The former reduces to the latter when $P_{s}$ is chosen as a Dirac distribution. We have the following lemma to upper bound the inverse regret under this mix-loss–based decomposition.
@@
Let $\{P_{s}\}_{s=1}^{\infty}$ be a stochastic process such that $P_{s}$ is a distribution over $\mathbb{R}^{d}$ and only relies on information collected until time s1s-1. Then, for any $\delta\in(0,1]$, we have
@@
Lemma 2 follows from Ville’s inequality (Ville, 1939), also known as “no-hypercompression” inequality (Grünwald, 2007, Chapter 3). It implies that term (a) in (9) is upper-bounded by $\log(1/\delta)$ when the mix loss is used as an intermediate term, significantly smaller than the $\mathcal{O}((\log t)^{3})$ bound in Zhang and Sugiyama (2023, Lemma 12) that employs the online logistic regression method (Foster et al., 2018). Notably, Lemma 2 allows flexible choices of $P_{s}$, enabling us to tailor $P_{s}$ to closely track the OMD estimator’s behavior. The formal statement is provided in Lemma 5 of Appendix A.
@@
$\sum_{s=1}^{t}m_{s}(P_{s})-\sum_{s=1}^{t}\ell_{s}(\theta_{s+1})\leq{}$ $\frac{1}{3\eta}\sum_{s=1}^{t}\lVert\theta_{s+1}-\theta_{s}\rVert^{2}_{H_{s}}+d(3\eta+\frac{1}{2})\ln\bigg(1+\frac{L_{\mu}t}{\lambda g(\tau)}\bigg),$
@@
More Technical Comparisons. Our analysis of the inverse regret is closely connected to prior work that bounds the cumulative negative log likelihood $L_{t}(\theta_{*})=\sum_{s=1}^{t}\ell_{s}(\theta_{*})$ using Ville’s inequality, a technique that traces back to Robbins (1970). Most existing approaches aim to bound $L_{t}(\theta_{*})$ with the loss of the MLE estimator, whereas our analysis focuses on the OMD estimator. This fundamental difference leads to a distinct construction of the intermediate term used in the decomposition.
In the context of GLB, our Lemma 2 resembles Lemma 3.3 of Lee et al. (2024), as both apply Ville’s inequality. However, their intermediate term $\mathbb{E}_{\theta\sim P_{t}}[L_{t}(\theta)]$ is built upon a distribution $P_{t}$ that is fixed across all individual functions $\{\ell_{s}\}_{s=1}^{t-1}$, making it naturally aligned with the MLE estimator but does not readily adapt to changing estimator. In contrast, our OMD estimator evolves with each individual function. Defining the intermediate term as the mix loss with time-varying $P_{s}$ thus provides the flexibility needed to track this changing comparator. Very recently, two concurrent works (Kirschner et al., 2025; Clerico et al., 2025) have also employed the mix loss to derive confidence sets. Our Lemma 2 corresponds to Proposition 2.1 of Clerico et al. (2025), and our mix loss aligns with the sequential likelihood mixing technique of Kirschner et al. (2025). The key distinction lies in the choice of estimator – MLE versus OMD – which leads to a fundamentally different specification of the mix loss and necessitates a tailored analysis. In their case, $P_{s}$ is defined as the outcome of the continuous Hedge algorithm, and the gap between the mix loss and the loss of any single comparator (including the MLE) is bounded using a standard regret guarantee. In contrast, our analysis must dynamically track the evolving OMD estimator, motivating our choice of $P_{s}$ as a Gaussian distribution with a time-varying mean $\{\theta_{s}\}_{s=1}^{t}$. Lemma 3 is then used to quantify the resulting loss gap.
@@ 5 Experiment
This section evaluates the proposed method on two representative GLB problems: logistic bandits ($\mu(z)=1/(1+e^{-z})$) with bounded rewards, and Poisson bandits ($\mu(z)=e^{z}$), which pose a distinct challenge as an unbounded GLB setting. We also conduct experiments on real data from the Covertype dataset (Blackard, 1998), with more detailed results provided in Appendix E.
Compared Methods. Four GLBs algorithms are compared, including GLM-UCB (Filippi et al., 2010), GLOC (Jun et al., 2017), RS-GLinCB (Sawarni et al., 2024), and OFUGLB (Lee et al., 2024). For logistic bandits, we further include two specialized algorithms: an MLE-based method with nearly optimal regret OFULog-r (Abeille et al., 2021), and a jointly efficient method ECOLog (Faury et al., 2022). We do not include OFUL-MLogB (Zhang and Sugiyama, 2023) since its confidence set is larger than that of ours, hence has a larger regret bound. More details of the baselines are provided in Appendix E. All experiments are conducted over 10 trials, and we report the average regret and running time.
@@
Results on Logistic Bandits. We conduct experiments under different configurations of $S$. The underlying parameter $\theta_{*}$ is sampled from a $d$-dimensional sphere with radius $S=\{3,5,7\}$, corresponding to $\kappa=21$, $137$, and $963$, respectively. Figure 1 reports the results. Among all methods, GLOC is the fastest but exhibits relatively large regret. OFUGLB attains the lowest regret due to its improved dependence on $\kappa$ and $S$, but as an MLE-based method, it incurs the highest computation cost. Our method strikes a favorable balance. Compared to OFUGLB, it achieves substantial cost savings with only a modest degradation in regret. Compared with ECOLog and RS-GLinCB, our method achieves comparable and even slightly better performance with improved computation cost. Moreover, it maintains a constant per-round cost across all regimes of $\kappa$, whereas the cost of RS-GLinCB increases with $\kappa$, as its rare-update strategy results in an update frequency that scales with $\kappa$.
Results on Poisson Bandits. We set the norm of the true parameter as $S\in\{3,5,7\}$, corresponding to $\kappa\approx 18$, $127$ and $882$. Poisson bandits have unbounded rewards, whereas GLM-UCB and RS-GLinCB require a predefined upper bound on the maximum reward as a parameter. We set it as 100, assuming rewards are effectively bounded with high probability. As shown in Figure 2, our method reduces the computational cost of OFUGLB by roughly 1000 times, with only a modest increase in regret.
@@ 6 Conclusion
This paper proposed a new method for the GLB problem that achieves a nearly optimal regret bound of $\mathcal{O}(\log T\sqrt{T/\kappa_{*}})$ with $\mathcal{O}(1)$ time and space complexities per round. Our approach builds on a novel analysis of the OMD estimator using the mix loss, enabling a tight confidence set construction for arm selection. A natural extension is to incorporate a warm-up strategy, as in prior work, to improve the dependence on $S$ and obtain $\kappa_{\mathcal{X}}$-based bounds. It also remains open whether geometry-aware bounds for GLBs can be achieved similar to those in logistic bandits (Abeille et al., 2021). Other directions include relaxing the self-concordance assumption toward weaker conditions (Liu et al., 2024), or improving $d$-dependence in the finite-arm setting (Jun et al., 2021; Mason et al., 2022).
Acknowledgments and Disclosure of Funding
Peng Zhao was supported by NSFC (62206125) and the Xiaomi Foundation. MS was supported by the Institute for AI and Beyond, UTokyo. The authors thank the reviewers for their valuable feedback and for bringing to our attention the recent works (Kirschner et al., 2025; Clerico et al., 2025).